# MODELING SKIERS FLOWS VIA WARDROPE EQUILIBRIUM IN CLOSED CAPACITATED NETWORKS

### v1

@@
Demyan Yarmoshik Moscow Institute of Physics and Technology 9, Institutskiy per., Dolgoprudny 141700, Russia yarmoshik.dv@phystech.edu
Igor Ignashin Moscow Institute of Physics and Technology 9, Institutskiy per., Dolgoprudny 141700, Russia ignashin.in@phystech.edu
Ekaterina Sikacheva Moscow Institute of Physics and Technology 9, Institutskiy per., Dolgoprudny 141700, Russia sikacheva.ev@phystech.edu
Alexander Gasnikov Moscow Institute of Physics and Technology 9, Institutskiy per., Dolgoprudny 141700, Russia Innopolis University 1, Universitetskaya Str., Innopolis, 420500, Russia Institute for Information Transmission Problems 19/1, Bolshoy Karetny per., Moscow, 127051, Russia gasnikov@yandex.ru
@@ 3.1 Equilibrium and utility
  • Its capacity $b_{u}$ measured in persons per unit of time. We assume that lift capacities $b\in\mathbb{R}^{|U|}_{+}$ are normalized by the total number of skiers in the system. Alike [2, 8, 9], we use a simple deterministic point queue model, so that the flow on the lift $f_{u}$ depends on the number of users in the queue $q_{u}$ and the input flow on the lift $f^{\text{in}}_{u}$ as
@@ 3.2 Optimization problem and competitive equilibrium
According to Wardrop’s principle, a transportation system, after sufficient time, reaches an equilibrium state in which the utility function is equal across all cycles. This competitive equilibrium corresponds to a stable distribution of traffic flows represented by a vector $\mathbf{n}^{*}\in\Delta^{C}$, where $\Delta^{C}$ is the unit simplex. The main objective is to determine this competitive equilibrium distribution. The related work by J.D. Brooks [17] investigates equilibria on cycles within bipartite graphs, providing foundational insights relevant to our study.
@@ 3.3 Queue network as nonlinear system of inequalities
Let $f^{C}\in\mathbb{R}^{|C|}_{+}$ and $f^{U}\in\mathbb{R}^{|U|}_{+}$ denote the auxiliary variables for flows on cycles and lifts respectively, and $\widehat{t}\in\mathbb{R}^{|C|}_{+}$ denote queue-free cycle passage times. Finally, let the matrix $\Theta$ encodes the relation between lifts and cycles: $\Theta_{cu}=1$ if and only if lift $u$ belongs to cycle $c$.
@@
$\widehat{\mathbf{t}}+\Theta\mathbf{t}$ $=\frac{\mathbf{n}}{\mathbf{f}^{C}}$ (3.4a)
$\Theta^{\top}\mathbf{f}^{C}$ $=\mathbf{f}^{U}$ (3.4b)
@@ 3.4 Toy networks
  • $t_{1}=0$, $t_{2}>0$. Using that the flow on the upper lift is equal to its capacity with (3.4a), we have two equations on $f^{C}_{1}$ and $t_{2}$: $f^{C}_{2}=b_{2}-f^{C}_{1}=\frac{n_{2}}{t_{2}+\widehat{t}_{2}}$, $f_{1}^{C}=\frac{n_{1}}{t_{2}+\widehat{t}_{1}}$. Substituting $f_{1}^{C}$, we obtain $b_{2}-\frac{n_{1}}{t_{2}+\widehat{t}_{1}}=\frac{n_{2}}{t_{2}+\widehat{t}_{2}}$. Since denominators are greater than zero, this is equivalent to a quadratic equation, whose only nonzero root (we prove uniqueness of waiting time values in the following subsection) is $t_{2}$.
The solution to the system can be found by considering solutions for all four cases and selecting the feasible one. This example shows that even such very simple network demonstrates complex behavior with a nonsmooth, nonlinear, nonconvex/nonconcave dependence of waiting times as functions of user distribution.
@@ 3.5 Computation of waiting times and cycle flows given a distribution of users
Having established the equilibrium conditions as a variational inequality, it is necessary to explicitly determine the waiting times in queues for any given distribution $\mathbf{n}$. This step is essential because the cycle utilities $\mathbf{\boldsymbol{\tau}}(\mathbf{n})$ depend explicitly on this quantity.
@@
$\min_{\mathbf{f}^{C}\geqslant 0}\quad\langle\widehat{\mathbf{t}},\mathbf{f}^{C}\rangle-\langle\mathbf{n},\ln\mathbf{f}^{C}\rangle\quad\text{subject to}\quad\Theta^{\top}\mathbf{f}^{C}\leqslant\mathbf{b},$ (3.5)
@@
$\mathscr{L}(\mathbf{f}^{C},\mathbf{t})=\langle\widehat{\mathbf{t}},\mathbf{f}^{C}\rangle-\langle\mathbf{n},\ln\mathbf{f}^{C}\rangle+\langle\mathbf{t},\Theta^{\top}\mathbf{f}^{C}-\mathbf{b}\rangle,$
@@
$\nabla_{\mathbf{f}^{C}}\mathscr{L}=\widehat{\mathbf{t}}-\frac{\mathbf{n}}{\mathbf{f}^{C}}+\Theta\mathbf{t}=\mathbf{0}\quad\Leftrightarrow\quad\widehat{\mathbf{t}}+\Theta\mathbf{t}=\frac{\mathbf{n}}{\mathbf{f}^{C}},$
@@
$\mathbf{t}\geqslant 0,\quad\Theta^{\top}\mathbf{f}^{C}\leqslant\mathbf{b},$
@@
$\mathbf{t}\odot(\Theta^{\top}\mathbf{f}^{C}-\mathbf{b})=\mathbf{0}.$
@@
$\max_{\mathbf{t}\geqslant 0}\min_{\mathbf{f}^{C}\geqslant 0}\quad\langle\widehat{\mathbf{t}}+\Theta\mathbf{t},\mathbf{f}^{C}\rangle-\langle\mathbf{n},\ln\mathbf{f}^{C}\rangle-\langle\mathbf{t},\mathbf{b}\rangle.$
Denoting $\mathbf{s}=\widehat{\mathbf{t}}+\Theta\mathbf{t}$ and finding that $\mathbf{f}^{C}(\mathbf{t})=\frac{\mathbf{n}}{\mathbf{s}}$ this equals to
@@
Since both primal (3.5) and dual (3.6) problems have strictly convex(concave) objectives, the solution $(\mathbf{f}^{C},\mathbf{t})$ corresponding to any fixed distribution $\mathbf{n}$ is unique. This establishes a bijection between the distribution of users $\mathbf{n}$ and the pair $(\mathbf{f}^{C},\mathbf{t})$ of cycle flows and waiting times.
Solving this system for $(\mathbf{f}^{C},\mathbf{t})$ using, e.g., an interior-point method provides the waiting times necessary for evaluation of the cycles utilities $\mathbf{\boldsymbol{\tau}}(\mathbf{n})$ in the variational inequality framework and allows to obtain the distribution of user flows in an approximate solution to the equilibrium problem.
@@ 4 Numerical experiments
We conducted experiments on a synthetic transportation network which is a larger variant of our toy network (Figure 1) consisting of five cycles and five lifts, see Figure 3. We used CVXPY package [20] for solving (3.5) to find the flows and waiting times for given user distribution $\mathbf{n}$, as described in Section 3.5. In our tests on the two-lifts toy network this implementation was always finding the solution to (3.4) with high accuracy.

### v2

@@
Demyan Yarmoshik
Moscow Institute of Physics and Technology
9, Institutskiy per., Dolgoprudny 141700, Russia
yarmoshik.dv@phystech.edu
Igor Ignashin
Moscow Institute of Physics and Technology
9, Institutskiy per., Dolgoprudny 141700, Russia
ignashin.in@phystech.edu
Ekaterina Sikacheva
Moscow Institute of Physics and Technology
9, Institutskiy per., Dolgoprudny 141700, Russia
sikacheva.ev@phystech.edu
Alexander Gasnikov
Moscow Institute of Physics and Technology
9, Institutskiy per., Dolgoprudny 141700, Russia
Innopolis University
1, Universitetskaya Str., Innopolis, 420500, Russia
Institute for Information Transmission Problems
19/1, Bolshoy Karetny per., Moscow, 127051, Russia
gasnikov@yandex.ru
@@ 3.1 Equilibrium and utility
  • Its capacity $b_{u}$ measured in persons per unit of time. We assume that lift capacities $b\in\mathbb{R}^{|U|}_{+}$ are normalized by the total number of skiers in the system. Similarly to [2, 8, 9], we use a simple deterministic point queue model, so that the flow on the lift $f_{u}$ depends on the number of users in the queue $q_{u}$ and the input flow on the lift $f^{\text{in}}_{u}$ as
@@ 3.2 Optimization problem and competitive equilibrium
According to Wardrop’s principle, a transportation system, after sufficient time, reaches an equilibrium state in which the utility function is equal across all cycles. This competitive equilibrium corresponds to a stable distribution of traffic flows represented by a vector $\mathbf{n}^{*}\in\Delta^{C}$, where $\Delta^{C}$ is the unit simplex. The main objective is to determine this competitive equilibrium distribution. The related work [16] investigates equilibria on cycles within bipartite graphs, providing foundational insights relevant to our study.
@@ 3.3 Queue network as nonlinear system of inequalities
Let $f^{C}\in\mathbb{R}^{|C|}_{+}$ and $f^{U}\in\mathbb{R}^{|U|}_{+}$ denote the auxiliary variables for flows on cycles and lifts respectively, and $\widehat{t}\in\mathbb{R}^{|C|}_{+}$ denote queue-free cycle passage times. Finally, let the matrix $\mathbf{\Theta}$ encode the relation between lifts and cycles: $\mathbf{\Theta}_{cu}=1$ if and only if lift $u$ belongs to cycle $c$.
@@
$\widehat{\mathbf{t}}+\mathbf{\Theta}\mathbf{t}$ $=\frac{\mathbf{n}}{\mathbf{f}^{C}}$ (3.4a)
$\mathbf{\Theta}^{\top}\mathbf{f}^{C}$ $=\mathbf{f}^{U}$ (3.4b)
@@ 3.4 Toy networks
  • $t_{1}=0$, $t_{2}>0$. Using that the flow on the upper lift is equal to its capacity with (3.4a), we have two equations on $f^{C}_{1}$ and $t_{2}$: $f^{C}_{2}=b_{2}-f^{C}_{1}=\frac{n_{2}}{t_{2}+\widehat{t}_{2}}$, $f_{1}^{C}=\frac{n_{1}}{t_{2}+\widehat{t}_{1}}$. Substituting $f_{1}^{C}$, we obtain $b_{2}-\frac{n_{1}}{t_{2}+\widehat{t}_{1}}=\frac{n_{2}}{t_{2}+\widehat{t}_{2}}$. Since denominators are greater than zero, this is equivalent to a quadratic equation, whose only nonzero root (we prove uniqueness of cycle waiting times in the following subsection) is $t_{2}$.
The solution to the system can be found by considering solutions for all four cases and selecting the feasible one. This example shows that even such very simple network demonstrates complex behavior with a nonsmooth, nonlinear, nonconvex/nonconcave dependence of cycle waiting times as functions of user distribution.
@@ 3.5 Computation of waiting times and cycle flows given a distribution of users
Having established the equilibrium conditions as a variational inequality, it is necessary to explicitly determine the cycle waiting times in queues for any given distribution $\mathbf{n}$. This step is essential because the cycle utilities $\mathbf{\boldsymbol{\tau}}(\mathbf{n})$ depend explicitly on this quantity.
@@
$\min_{\mathbf{f}^{C}\geqslant 0}\quad\langle\widehat{\mathbf{t}},\mathbf{f}^{C}\rangle-\langle\mathbf{n},\ln\mathbf{f}^{C}\rangle\quad\text{subject to}\quad\mathbf{\Theta}^{\top}\mathbf{f}^{C}\leqslant\mathbf{b},$ (3.5)
@@
$\mathscr{L}(\mathbf{f}^{C},\mathbf{t})=\langle\widehat{\mathbf{t}},\mathbf{f}^{C}\rangle-\langle\mathbf{n},\ln\mathbf{f}^{C}\rangle+\langle\mathbf{t},\mathbf{\Theta}^{\top}\mathbf{f}^{C}-\mathbf{b}\rangle,$
@@
$\nabla_{\mathbf{f}^{C}}\mathscr{L}=\widehat{\mathbf{t}}-\frac{\mathbf{n}}{\mathbf{f}^{C}}+\mathbf{\Theta}\mathbf{t}=\mathbf{0}\quad\Leftrightarrow\quad\widehat{\mathbf{t}}+\mathbf{\Theta}\mathbf{t}=\frac{\mathbf{n}}{\mathbf{f}^{C}},$
@@
$\mathbf{t}\geqslant 0,\quad\mathbf{\Theta}^{\top}\mathbf{f}^{C}\leqslant\mathbf{b},$
@@
$\mathbf{t}\odot(\mathbf{\Theta}^{\top}\mathbf{f}^{C}-\mathbf{b})=\mathbf{0}.$
@@
$\max_{\mathbf{t}\geqslant 0}\min_{\mathbf{f}^{C}\geqslant 0}\quad\langle\widehat{\mathbf{t}}+\mathbf{\Theta}\mathbf{t},\mathbf{f}^{C}\rangle-\langle\mathbf{n},\ln\mathbf{f}^{C}\rangle-\langle\mathbf{t},\mathbf{b}\rangle.$
Denoting $\mathbf{s}=\widehat{\mathbf{t}}+\mathbf{\Theta}\mathbf{t}$ and finding that $\mathbf{f}^{C}(\mathbf{t})=\frac{\mathbf{n}}{\mathbf{s}}$ this equals to
@@
Since primal (3.5) problem has strictly convex objective, its solution $\mathbf{f}^{C}$ corresponding to any fixed distribution $\mathbf{n}$ is unique. A similar statement holds for the dual problem (3.6): its objective is strictly concave on $\ker^{\bot}\mathbf{\Theta}$. Since changing $\mathbf{t}$ along $\ker\mathbf{\Theta}$ does not affect cycle waiting times $\mathbf{\Theta}\mathbf{t}$, this establishes a bijection between the distribution of users $\mathbf{n}$ and the pair $(\mathbf{f}^{C},\mathbf{\Theta}\mathbf{t})$ of cycle flows and cycle waiting times.
Solving this system for $(\mathbf{f}^{C},\mathbf{\Theta}\mathbf{t})$ using, e.g., an interior-point method provides the cycle waiting times necessary for evaluation of the cycles utilities $\mathbf{\boldsymbol{\tau}}(\mathbf{n})$ in the variational inequality framework and allows to obtain the distribution of user flows in an approximate solution to the equilibrium problem.
@@ 4 Numerical experiments
We conducted experiments on a synthetic transportation network which is a larger variant of our toy network (Figure 1) consisting of five cycles and five lifts, see Figure 3. We used CVXPY package [19] for solving (3.5) to find the cycle flows and cycle waiting times for given user distribution $\mathbf{n}$, as described in Section 3.5. In our tests on the two-lifts toy network this implementation was always finding the solution to (3.4) with high accuracy.